# EarthGPT-X: A Spatial MLLM for Multi-level Multi-Source Remote Sensing Imagery Understanding with Visual Prompting

### v2

@@ Abstract
Recent advances in natural-domain multi-modal large language models (MLLMs) have demonstrated effective spatial reasoning through visual and textual prompting. However, their direct transfer to remote sensing (RS) is hindered by heterogeneous sensing physics, diverse modalities, and unique spatial scales. Existing RS MLLMs are mainly limited to optical imagery and plain language interaction, preventing flexible and scalable real-world applications. In this article, EarthGPT-X is proposed, the first flexible spatial MLLM that unifies multi-source RS imagery comprehension and accomplishes both coarse-grained and fine-grained visual tasks under diverse visual prompts in a single framework. Distinct from prior models, EarthGPT-X introduces: 1) a dual-prompt mechanism combining text instructions with various visual prompts (i.e., point, box, and free-form) to mimic the versatility of referring in human life; 2) a comprehensive multi-source multi-level prompting dataset, the model advances beyond holistic image understanding to support hierarchical spatial reasoning, including scene-level understanding and fine-grained object attributes and relational analysis; 3) a cross-domain one-stage fusion training strategy, enabling efficient and consistent alignment across modalities and tasks. Extensive experiments demonstrate that EarthGPT-X substantially outperforms prior nature and RS MLLMs, establishing the first framework capable of multi-source, multi-task, and multi-level interpretation using visual prompting in RS scenarios. The code and dataset are available at https://github.com/wivizhang/EarthGPT-X.
@@ I Introduction
Typical spatial understanding tasks in multi-modal learning require aligning image regions with text semantics[1]. Beyond holistic description, such tasks involve accurately interpreting specific regions or objects based on user instructions[2, 3]. Most recently, enabling multi-modal large language models (MLLMs) with comprehensive spatial understanding has become the focus of research in natural scenarios[4, 5, 6]. However, these models struggle with the remote sensing (RS) imagery, which features multi-source, multi-viewpoint, and scale variations due to different imaging conditions[7]. Although existing RS MLLMs can handle multiple RS visual tasks[8, 9], they primarily focus on holistic image understanding, lacking spatial perception. They solely rely on text instructions, which can lead to ambiguity and limited context. It is indispensable to develop a spatial-aware MLLM to interpret RS imagery under flexible interactions.
Notably, in the human world, people can achieve an all-around understanding of visual information, allowing for both zoom-in fine-grained analyses and zoom-out coarse-grained overviews[10]. Inspired by this, visual prompting models advance the field by injecting visual prompts beyond text instructions to enable multi-grained visual reasoning and enhance spatial perception[11]. Representative works include the promptable segmentation model Segment Anything (SAM) [12], which enables class-agnostic delineation from visual or text prompts and shows strong generalization as a versatile backbone. The region-of-interest model GPT4RoI [4] integrates large language models (LLMs) for localized reasoning, while Ferret [1] focuses on fine-grained grounding, which can accept diverse region inputs, such as points and bounding boxes. Additionally, Wu et al. [13] provide a review of visual prompting approaches, including visual prompting, prompt generation, and prompt learning. However, designed for natural images, these models cannot be directly adapted to the heterogeneous RS domain due to fundamental differences in sensing physics, modalities, and scales.
Figure 1: Overview of the proposed framework for flexible and comprehensive spatial understanding in RS imagery: (a) Tasks: a wide range of multi-granularity visual tasks; (b) Model: EarthGPT-X integrates an image encoder, a visual prompt encoder, a text tokenizer, an LLM, and a pixel perception module; (c) Data: the training dataset M-RSVP is constructed from multi-source RS data (optical, SAR, infrared) and further enriched with GPT-4V–based annotations.
@@
To overcome these limitations, we propose EarthGPT-X, a spatial MLLM built upon visual prompt learning, tailored for multi-source RS imagery, supporting diverse visual prompt types (e.g., points, boxes, circles, and scribbles), and enabling comprehensive interpretation from overall scene understanding to fine-grained regional analysis, thereby offering more flexible and scalable visual-language interaction than prior approaches. As depicted in Fig. 1, the EarthGPT-X architecture consists of three interconnected components: multi-grained RS image interpretation tasks, a visual prompting model, and a construction pipeline of training datasets. Specifically, EarthGPT-X not only allows for zoom-in fine-grained reasoning (e.g., free-form region captioning, referring object classification, and object inter-relationship analyses), but also supports zoom-out coarse-grained tasks (e.g., image-level captioning, scene classification). Secondly, the key of the model is to realize multi-modal content integration, including images, visual prompts, and text instructions. The multi-modal signals are then fed into the LLMs for a one-stage all-in-one training. Additionally, a pixel perception module is integrated to extend the model with object segmentation masks. Most importantly, the training dataset named M-RSVP is constructed, transforming from existing multi-sensor datasets, which are uniformly reorganized and relabeled as image-visual prompts-text triples. In addition, a portion of the data re-annotated using GPT-4V[16] to generate longer and more detailed descriptions. A side-by-side comparison of EarthGPT-X with recent RS MLLMs is presented in Tab. I, highlighting its advancement in modality coverage, dataset quality, prompting flexibility, and fine-grained multi-task reasoning.
To further validate EarthGPT-X’s comprehensive interpretation ability on multi-source imagery, we compare it with recent state-of-the-art (SOTA) natural and RS MLLMs. The results indicate that EarthGPT-X achieves accurate multi-source reasoning and provides more detailed RS image understanding. Overall, EarthGPT-X not only attains competitive performance on vision–language benchmarks, but also demonstrates remarkable advantages in complex object-level reasoning tasks.
@@
  • M-RSVP: First Multi-source Visual Prompting Dataset. The first multi-source visual prompting dataset for RS, titled M-RSVP, consisting of over 650k image-visual prompt-text triples, is constructed. The constructed M-RSVP bridges the gap of multi-source visual prompting data in the RS domain, enabling multi-granularity interpretation across diverse scenes and sensors.
  • Multi-domain Multi-modal Multi-level Interaction. A hybrid signal mutual understanding mechanism is designed to strengthen the interplay between dense image features, sparse visual prompts, and text instructions, thereby improving the interpretation of referring regions within holistic images. Moreover, our cross-domain one-stage fusion training significantly streamlines the training procedure of MLLMs, while ensuring seamless integration of multi-source knowledge.
  • Superior Performance. Extensive experiments benchmark EarthGPT-X against SOTA natural and RS MLLMs across diverse multi-level visual tasks and multi-source imagery. Furthermore, EarthGPT-X also excels in complex reasoning tasks, such as object inter-relationship analysis. These advances collectively represent a significant milestone toward developing versatile and flexible MLLMs for RS scenarios.
@@ II-C Prompting Learning
Prompt learning has become a key research direction in NLP [24], with works such as AutoPrompt [50] and CoOp [51] automating prompt design for language and vision–language models. Subsequently, the idea evolved into visual prompting [13], marking a broader shift from language-based to multi-modal prompts. A milestone in this line is the Segment Anything Model (SAM) [12], which enables class-agnostic segmentation via point, box, or text prompts and demonstrates strong generalization across diverse domains. GPT4RoI [4] and RegionBlip [5] leverage spatial boxes and combine language with region-of-interest input for regional recognition. Colorful Prompting Tuning (CPT) [52] uses color-based prompts to enhance MLLMs’ performance. Osprey [3] goes further by incorporating fine-grained mask regions into language instructions for pixel-level understanding. ViP-LLava[47] supports arbitrary visual cues like scribbles and arrows to interpret images flexibly. Wu et al. [13] provide a comprehensive survey of visual prompting approaches, covering visual prompting, prompt generation, and prompt learning. These models are primarily designed for the natural image domain and fail to generalize to the RS multi-source data. Notably, EarthMarker [15] is a visual prompting framework tailored for the RS data, but it remains confined to optical imagery and does not support flexible prompt types. These gaps highlight the necessity of a specialized model that can support multi-source imagery and flexible interaction in RS scenarios, motivating the development of the proposed model.
@@ III-A Multi-modal Content Integration
The multi-modal input stream of the EarthGPT-X is characterized by its tripartite structure, comprising an image global and local perception module, a spatial-aware module, and a text instruction processing module. The multi-modal content encoding and integration are detailed as follows.
@@ III-A4 Hybrid Signals Mutual Understanding
$[V\text{-}P]_{\mathrm{hybrid}}$ $=\mathrm{Self\text{-}attention}~(V_{\mathrm{ca}}+R_{\mathrm{ff}}),$ (7)
@@ III-C Integrated with Pixel Perception Module
$M=\mathcal{P}\Big(\mathcal{R}_{\mathrm{lp}}(L_{\mathrm{pix}}),\mathcal{G}(I)\Big).$ (13)
@@ IV-C GPT4V-assisted Data Generation
To enable high-quality annotations with GPT-4V [76], we followed the design principles of [77] and developed a set of rigorous and standardized prompt templates, as illustrated in Fig. 4. These templates are tailored to distinct tasks: brief captioning, detailed captioning, and inter-relationship analysis. Notably, category names and ground-truth bounding boxes are explicitly embedded in the prompts, ensuring the correctness of the generated annotations by guiding GPT-4V to accurately associate text with visual regions. Beyond correctness, the structured templates also guarantee annotation consistency across large-scale data, avoiding ambiguity and hallucination. To further strengthen the coupling between visual prompts and textual outputs, we adopt Set-of-Marks (SoM) prompting [78], which highlights multiple target objects simultaneously and reinforces spatial-textual grounding. This design enables GPT-4V to produce coherent and complete descriptions even in cluttered scenes. In conclusion, the integration of diverse multi-domain datasets through this rigorous pipeline allows joint training of EarthGPT-X on standardized supervision, enhancing its robustness across RS modalities and improving its generalization to complex real-world scenarios. Additionally, to further ensure reliability, we conduct RS expert inspection on sampled annotations, validating semantic accuracy, coherence, and completeness. This human-in-the-loop verification step complements automatic generation and safeguards the overall quality of the constructed dataset.
@@ V-A2 Evaluation Metrics
We adopt three categories of evaluation metrics tailored to different visual tasks in our benchmark. (i) Region and Image Captioning tasks: We employ standard language generation metrics, including BLEU [81], METEOR [82], ROUGE [83], CIDEr [84], and SPICE [85]. (ii) Scene Classification: We report classification accuracy, which directly reflects the model’s ability to recognize high-level scene semantics in RS images. (iii) Referring Object Classification: We employ two semantic-aware metrics: Semantic Intersection over Union (S-IOU) and Semantic Similarity (SS) [86, 3, 77]. SS measures the similarity of predicted and ground-truth labels in a semantic embedding space, while S-IOU captures the semantic overlap between them by computing the intersection-over-union in the embedding space. Those metrics provide a comprehensive assessment of the quality of the identified object information and generated captions. Consider the referring expression “A airplane is on the right of a small vehicle”, where the ground-truth target labels are airplane and vehicle. Suppose the model predicts aircraft and car, both outputs achieve high SS since they are semantically close to the labels in the embedding space, and also high S-IoU because of strong token-level overlap. In contrast, if the predictions are ship and building, SS drops due to weaker semantic similarity, while S-IoU becomes very low as the key object words differ significantly from the ground-truth labels.
@@
Models UCMerced Acc AID Acc
Qwen-VL [87] 62.90 52.60
MiniGPTV2 [88] 4.76 12.90
LLaVa-1.5 [29] 68.00 51.00
Sphinx [89] 62.76 58.20
InternVL [90] 58.23 60.40
GeoChat [9] 84.43 72.03
EarthMarker [15] 86.52 77.97
@@
Models Point Box (Coor) Free-form
@@
InternVL2.5 [38] 54.54 37.91
InternVL3 [91] 50.84 26.99
GeoChat [9] 79.59 68.80
Sphinx [89] 93.72 89.37
EarthGPT [8] 94.64 90.16
@@
Sphinx -V [77] 86.21 78.78 89.07 81.62 88.46 79.97
Vip-LLava-7b [18] 68.32 51.98 72.56 55.94 71.24 53.99
Vip-LLava-13b [18] 70.02 57.61 74.51 60.53 74.05 59.83
EarthMarker [15] 95.96 93.49 98.37 97.24
EarthGPT-X (Ours) 96.02 94.12 98.60 97.83 98.30 96.56
Notes: Since vanilla MLLMs lack visual prompt support, textual coordinates are treated as the referring regions. ‘—’ means no such capability.
@@
Models Formats $S$ S-IoU
InternVL2.5 [38] Coor 37.05 7.69
InternVL3 [91] Coor 32.72 2.89
Sphinx-V [77] Box 34.76 2.03
@@
Models BLEU1 BLEU2 BLEU3 BLEU4 METEOR ROUGE-L CIDEr(0-5) SPICE
Expert Models
CSMLF [92] 77.0 64.9 53.2 47.1 32.0 57.8 106.5 26.5
Qu, et. al [60] 72.5 60.3 51.8 45.5 33.6 59.1 117.9 27.6
Attention (soft) [59] 73.1 60.9 52.5 46.2 33.9 59.9 113.6 28.5
Attention (hard) [59] 73.3 61.0 52.7 46.4 34.0 60.0 110.3 28.4
FC-Att [93] 73.6 61.5 53.2 46.9 33.8 60.0 123.1 28.3
SM-Att [93] 73.9 61.7 53.2 46.8 33.0 59.3 123.6 27.6
MLCA-Net [94] 74.5 62.4 54.1 47.8 33.7 60.1 126.4 28.5
@@
Models Format BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE CIDEr SPICE
@@
Qwen-VL [9] Coor 12.78 6.47 2.77 1.24 5.97 17.06 13.15 6.60
InternVL2.5 [38] Coor 9.87 4.18 1.69 0.52 5.71 10.55 9.22 8.28
InternVL3 [91] Coor 15.42 8.91 6.34 4.38 18.64 14.28 13.27 14.87
GeoChat [87] Coor 15.02 8.49 5.21 3.29 9.24 25.65 42.86 16.10
Sphinx [89] Coor 22.79 15.27 10.97 7.84 12.95 30.28 105.85 20.18
EarthGPT [8] Coor 30.28 19.99 13.58 8.73 15.29 31.52 88.92 21.05
@@
Sphinx-V [77] Box 23.35 14.84 10.37 7.16 14.10 32.08 81.79 22.46
Vip-LLava-7b [18] Box 14.67 7.29 4.06 2.25 8.20 20.17 23.47 12.79
Vip-LLava-13b [18] Box 16.39 8.40 4.77 2.80 8.47 21.36 25.85 12.41
EarthGPT-X (Ours) Point 52.22 48.13 45.15 42.88 35.01 64.31 477.40 63.14
EarthGPT-X (Ours) Box 64.55 59.76 56.10 53.11 39.22 66.64 487.44 64.20
EarthGPT-X (Ours) Free-form 63.25 58.16 54.89 52.01 38.35 65.01 482.26 63.23
@@
Models SAR-Ship (Supervised) HIT UAV (Zero-Shot)
@@
GeoChat[9] 57.15 57.15 52.20 59.85 59.85 51.31
EarthDial[23] 63.10 63.10 54.83 61.83 61.83 52.80
@@
V-B Scene Classification

### v3

@@ Abstract
Recent advances in natural-domain multi-modal large language models (MLLMs) have demonstrated effective spatial reasoning through visual and textual prompting. However, their direct transfer to remote sensing (RS) is hindered by heterogeneous sensing physics, diverse modalities, and unique spatial scales. Existing RS MLLMs are mainly limited to optical imagery and plain language interaction, preventing flexible and scalable real-world applications. In this article, EarthGPT-X is proposed, the first flexible spatial MLLM that unifies multi-source RS imagery comprehension and accomplishes both coarse-grained and fine-grained visual tasks under diverse visual prompts in a single framework. Distinct from prior models, EarthGPT-X introduces: 1) a dual-prompt mechanism combining text instructions with various visual prompts (i.e., point, box, and free-form) to mimic the versatility of referring in human life; 2) a comprehensive multi-source multi-level prompting dataset, the model advances beyond holistic image understanding to support hierarchical spatial reasoning, including scene-level understanding and fine-grained object attributes and relational analysis; 3) a cross-domain one-stage fusion training strategy, enabling efficient and consistent alignment across modalities and tasks. Extensive experiments demonstrate that EarthGPT-X substantially outperforms prior natural and RS MLLMs, establishing the first framework capable of multi-source, multi-task, and multi-level interpretation using visual prompting in RS scenarios. The code and dataset are available at https://github.com/wivizhang/EarthGPT-X.
@@ I Introduction
Typical spatial understanding tasks in multi-modal learning require aligning image regions with text semantics[1]. Beyond holistic analysis, spatial tasks involve accurately interpreting specific regions or objects based on user instructions[2, 3]. Most recently, enabling multi-modal large language models (MLLMs) with comprehensive spatial understanding has become the focus of research in natural scenarios[4, 5, 6]. However, these models struggle with the remote sensing (RS) imagery, which features multi-source, multi-viewpoint, and scale variations due to different imaging conditions[2]. Although existing RS MLLMs can handle multiple RS visual tasks[7, 8], they primarily focus on holistic image understanding, lacking spatial perception. They solely rely on text instructions, which can lead to ambiguity and limited context. It is indispensable to develop a spatial-aware MLLM to interpret RS imagery under flexible interactions.
Notably, in the human world, people can achieve an all-around understanding of visual information, allowing for both zoom-in fine-grained analyses and zoom-out coarse-grained overviews[9]. Inspired by this, visual prompting models advance the field by injecting visual prompts beyond text instructions to enable multi-grained visual reasoning and enhance spatial perception[10]. Representative works include the promptable segmentation model Segment Anything (SAM) [11], which enables class-agnostic delineation from visual or text prompts and shows strong generalization as a versatile backbone. The region-of-interest model GPT4RoI [4] integrates large language models (LLMs) for localized reasoning, while Ferret [1] focuses on fine-grained grounding, which can accept diverse region inputs, such as points and bounding boxes. Additionally, Wu et al. [12] provide a review of visual prompting approaches, including prompt generation and prompt learning. However, designed for natural images, these models cannot be directly adapted to the heterogeneous RS domain due to fundamental differences in sensing physics, modalities, and spatial resolutions.
Figure 1: Overview of the proposed framework for flexible and comprehensive spatial understanding in RS imagery: (a) Tasks: a wide range of multi-level visual reasoning tasks; (b) Model: EarthGPT-X integrates an image encoder, a visual prompt encoder, a text tokenizer, an LLM, and a pixel perception module; (c) Data: the training dataset M-RSVP is constructed from multi-source RS data (i.e., optical, synthetic aperture radar (SAR), infrared) and further enriched with GPT-4V–based annotations.
@@
To overcome these limitations, we propose EarthGPT-X, a spatial MLLM built upon visual prompt learning and tailored for multi-source RS imagery. It supports arbitrary visual prompt types (e.g., points, boxes, circles, and scribbles) and enables comprehensive interpretation from overall scene understanding to fine-grained regional analysis, thereby providing more flexible and scalable visual–language interaction than prior approaches. As depicted in Fig. 1, the EarthGPT-X architecture consists of three interconnected components: multi-level RS image understanding tasks, a visual prompting model, and a construction pipeline of training dataset. Specifically, EarthGPT-X not only allows for zoom-in fine-grained reasoning (e.g., free-form region captioning, referring object classification, and object inter-relationship analysis), but also supports zoom-out coarse-grained tasks (e.g., image-level captioning, scene classification). Secondly, the key of the model is to realize multi-modal content integration, including images, visual prompts, and text instructions. The multi-modal signals are then fed into the LLMs for a one-stage all-in-one training. Additionally, a pixel perception module is integrated to extend the model with object segmentation masks. Most importantly, the training dataset named M-RSVP is constructed, transforming from existing multi-sensor datasets, which are uniformly reorganized and relabeled as image-visual prompts-text triples. In addition, a portion of the data re-annotated using GPT-4V[15] to generate longer and more detailed descriptions. A side-by-side comparison of EarthGPT-X with recent RS MLLMs is presented in Tab. I, highlighting its advancement in modality coverage, dataset quality, prompting flexibility, and fine-grained multi-task reasoning.
To further validate EarthGPT-X’s comprehensive interpretation ability on multi-source imagery, we compare it with recent state-of-the-art (SOTA) natural and RS MLLMs. The results indicate that EarthGPT-X achieves accurate multi-source reasoning and provides more detailed RS image understanding. Overall, EarthGPT-X not only attains competitive performance on visual–language benchmarks, but also demonstrates remarkable advantages in complex object-level reasoning tasks.
@@
  • M-RSVP: First Multi-source Visual Prompting Dataset. The first multi-source visual prompting dataset for RS, titled M-RSVP, consisting of over 650k image-visual prompt-text triples, is constructed. The constructed M-RSVP bridges the gap of multi-source visual prompting data in the RS domain, enabling multi-granularity image interpretation across diverse scenes and sensors.
  • Multi-domain Multi-modal Multi-level Interaction. A hybrid signal mutual understanding mechanism is designed to strengthen the interplay between dense image features, sparse visual prompts, and text instructions, thereby improving the interpretation of referring regions within holistic images. Moreover, our cross-domain one-stage fusion training significantly streamlines the training procedure of MLLMs, while ensuring seamless integration of multi-modal knowledge.
  • Superior Performance. Extensive experiments benchmark EarthGPT-X against SOTA natural and RS MLLMs across diverse multi-level visual tasks and multi-source imagery, outperforming existing models. Furthermore, EarthGPT-X also excels in complex reasoning tasks, such as object inter-relationship analysis. These advances collectively represent a significant milestone toward developing versatile and flexible MLLMs for RS scenarios.
@@ II-C Prompting Learning
Prompt learning has become a key research direction in NLP [23], with works such as AutoPrompt [49] and CoOp [50] automating prompt design for language and visual–language models. Subsequently, the idea evolved into visual prompting [12], marking a broader shift from language-based to multi-modal prompts. A milestone in this line is the SAM [11], which enables class-agnostic segmentation via point, box, or text prompts and demonstrates strong generalization across diverse domains. GPT4RoI [4] and RegionBlip [5] leverage spatial boxes and combine language with region-of-interest input for regional recognition. Colorful Prompting Tuning (CPT) [51] uses color-based prompts to enhance MLLMs’ performance. Osprey [3] goes further by incorporating fine-grained mask regions into language instructions for pixel-level understanding. ViP-LLava[46] supports arbitrary visual cues like scribbles and arrows to interpret images flexibly. Wu et al. [12] provide a comprehensive survey of visual prompting approaches, covering visual prompting, prompt generation, and prompt learning. These models are primarily designed for the natural image domain and fail to generalize to the RS multi-source data. Notably, EarthMarker [14] is a visual prompting framework tailored for the RS data, but it remains confined to optical imagery and does not support flexible prompt types. These gaps highlight the necessity of a specialized model that can support multi-source imagery and flexible interaction in RS scenarios, motivating the development of the proposed model.
@@ III-A Multi-modal Content Integration
The multi-modal input stream of the EarthGPT-X is characterized by its tripartite structure. The multi-modal content encoding and integration are detailed as follows.
@@ III-A4 Hybrid Signals Mutual Understanding
$[V\text{-}P]_{\mathrm{hybrid}}$ $=\mathrm{Self\text{-}attention}~(H_{\mathrm{ca}}+E_{\mathrm{ff}}),$ (7)
@@ III-C Integrated with Pixel Perception Module
$M=\mathcal{P}\Big(\mathcal{R}_{\mathrm{lp\text{-}p}}(L_{\mathrm{pix}}),\mathcal{G}(I)\Big).$ (13)
@@ IV-C GPT4V-assisted Data Generation
To enable high-quality annotations with GPT-4V [75], we follows the design principles of [76] and developed a set of rigorous and standardized prompt templates, as illustrated in Fig. 4. These templates are tailored to distinct tasks: brief captioning, detailed captioning, and inter-relationship analysis. Notably, category names and ground-truth bounding boxes are explicitly embedded in the prompts, ensuring the correctness of the generated annotations by guiding GPT-4V to accurately associate text with visual regions. Beyond correctness, the structured templates also guarantee annotation consistency across large-scale data, avoiding ambiguity and hallucination. To further strengthen the coupling between visual prompts and textual outputs, we adopt Set-of-Marks (SoM) prompting [77], which highlights multiple target objects simultaneously and reinforces spatial-textual grounding. This design enables GPT-4V to produce coherent and complete descriptions even in cluttered scenes. In conclusion, the integration of diverse multi-domain datasets through this rigorous pipeline allows joint training of EarthGPT-X on standardized supervision, enhancing its robustness across RS modalities and improving its generalization to complex real-world scenarios. Additionally, to further ensure reliability, we conduct RS expert inspection on sampled annotations, validating semantic accuracy, coherence, and completeness. This human-in-the-loop verification step complements automatic generation and safeguards the overall quality of the constructed dataset.
@@ V-A2 Evaluation Metrics
We adopt three categories of evaluation metrics tailored to different visual tasks in our benchmark. (i) Region and Image Captioning Tasks: We employ standard language generation metrics, including BLEU [80], METEOR [81], ROUGE [82], CIDEr [83], and SPICE [84]. (ii) Scene Classification: We report classification accuracy, which directly reflects the model’s ability to recognize high-level scene semantics in RS images. (iii) Referring Object Classification: We employ two semantic-aware metrics: Semantic Intersection over Union (S-IOU) and Semantic Similarity (SS) [85, 3, 76]. SS measures the similarity of predicted and ground-truth labels in a semantic embedding space, while S-IOU captures the semantic overlap between them by computing the intersection-over-union in the embedding space. These metrics provide a comprehensive assessment of the quality of the identified object information and generated captions. Consider the referring expression “An airplane is on the right of a small vehicle”, where the ground-truth target labels are airplane and vehicle. Suppose the model predicts aircraft and car, both outputs achieve high SS since they are semantically close to the labels in the embedding space, and also high S-IoU because of strong token-level overlap. In contrast, if the predictions are ship and building, SS drops due to weaker semantic similarity, while S-IoU becomes very low as the key object words differ significantly from the ground-truth labels.
V-B Scene Classification
@@
Models Publications UCMerced Acc AID Acc
Qwen-VL [88] arXiv 2023 62.90 52.60
MiniGPTV2 [89] arXiv 2023 4.76 12.90
LLaVa-1.5 [28] NeurIPS 2023 68.00 51.00
Sphinx [90] ECCV 2024 62.76 58.20
InternVL [91] CVPR 2024 58.23 60.40
GeoChat [8] CVPR 2024 84.43 72.03
EarthMarker [14] TGRS 2025 86.52 77.97
@@
Models Publications Point Box (Coor) Free-form
@@
InternVL2.5 [37] arXiv 2024 54.54 37.91
InternVL3 [92] arXiv 2025 50.84 26.99
GeoChat [8] CVPR 2024 79.59 68.80
Sphinx [90] ECCV 2024 93.72 89.37
EarthGPT [7] TGRS 2024 94.64 90.16
@@
Vip-LLava-7b [17] CVPR 2024 68.32 51.98 72.56 55.94 71.24 53.99
Vip-LLava-13b [17] CVPR 2024 70.02 57.61 74.51 60.53 74.05 59.83
Sphinx -V [76] ICLR 2025 86.21 78.78 89.07 81.62 88.46 79.97
EarthMarker [14] TGRS 2025 95.96 93.49 98.37 97.24
EarthGPT-X (Ours) 96.33 94.42 98.79 98.03 98.71 96.72
Notes: Since vanilla MLLMs lack visual prompt support, textual coordinates (abbr. as Coor) are treated as the referring regions. ‘—’ means no such capability.
The underline denotes the best performance achieved under box prompts.
@@
Models Publications Formats $S$ S-IoU
InternVL2.5 [37] arXiv 2024 Coor 37.05 7.69
InternVL3 [92] arXiv 2025 Coor 32.72 2.89
Sphinx-V [76] ICLR 2025 Box 34.76 2.03
@@
Models Publications BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE -L CIDEr(0-5) SPICE
RS Expert Models
CSMLF [93] GRSL 2019 77.0 64.9 53.2 47.1 32.0 57.8 106.5 26.5
Qu et al. [59] CITS 2016 72.5 60.3 51.8 45.5 33.6 59.1 117.9 27.6
Attention (soft) [58] TGRS 2017 73.1 60.9 52.5 46.2 33.9 59.9 113.6 28.5
Attention (hard) [58] TGRS 2017 73.3 61.0 52.7 46.4 34.0 60.0 110.3 28.4
FC-Att [94] TGRS 2019 73.6 61.5 53.2 46.9 33.8 60.0 123.1 28.3
SM-Att [94] TGRS 2019 73.9 61.7 53.2 46.8 33.0 59.3 123.6 27.6
MLCA-Net [95] TGRS 2022 74.5 62.4 54.1 47.8 33.7 60.1 126.4 28.5
Capformer [96] IGARSS 2022 89.0 81.1 74.5 69.2 45.4 78.4 201.9
Aware-Transformer [97] CGI 2023 91.5 85.2 79.6 75.0 47.7 81.1 214.7
RS MLLMs
RS-CapRet [98] arXiv 2024 87.1 78.7 71.7 65.6 43.6 77.6 192.9 31.1
EarthGPT [7] TGRS 2024 87.1 78.7 71.6 65.5 44.5 78.2 192.6 32.2
EarthMarker [14] TGRS 2025 84.4 73.1 62.9 54.3 37.5 70.0 162.9 26.8
@@
Models Publications Format BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE CIDEr SPICE
@@
Qwen-VL [88] arXiv 2023 Coor 12.78 6.47 2.77 1.24 5.97 17.06 13.15 6.60
InternVL2.5 [37] arXiv 2024 Coor 9.87 4.18 1.69 0.52 5.71 10.55 9.22 8.28
InternVL3 [92] arXiv 2025 Coor 15.42 8.91 6.34 4.38 18.64 14.28 13.27 14.87
GeoChat [8] CVPR 2024 Coor 15.02 8.49 5.21 3.29 9.24 25.65 42.86 16.10
Sphinx [90] ECCV 2024 Box 22.79 15.27 10.97 7.84 12.95 30.28 105.85 20.18
EarthGPT [7] TGRS 2024 Box 30.28 19.99 13.58 8.73 15.29 31.52 88.92 21.05
@@
Sphinx-V [76] ICLR 2025 Box 23.35 14.84 10.37 7.16 14.10 32.08 81.79 22.46
Vip-LLava-7b [17] CVPR 2024 Box 14.67 7.29 4.06 2.25 8.20 20.17 23.47 12.79
Vip-LLava-13b [17] CVPR 2024 Box 16.39 8.40 4.77 2.80 8.47 21.36 25.85 12.41
GLaMM [99] CVPR 2024 Box 12.50 6.06 3.25 1.74 7.98 22.01 23.85 12.40
EarthGPT-X (Ours) Point 52.57 48.14 45.72 43.38 35.28 64.63 479.92 63.47
EarthGPT-X (Ours) Box 64.83 59.91 56.28 53.64 39.61 66.87 488.76 64.32
EarthGPT-X (Ours) Free-form 63.38 58.62 55.12 52.41 38.72 65.43 483.63 63.58
@@
Models Publications SAR-Ship (Supervised) HIT UAV (Zero-Shot)
@@
GeoChat [8] CVPR 2024 57.15 57.15 52.20 59.85 59.85 51.31
EarthDial [22] CVPR 2025 63.10 63.10 54.83 61.83 61.83 52.80
@@
Notes: Results for region-level tasks are reported under box-prompt settings.